OpenStudy (frostarmy):

Will Fan and give Medal What is the linear function equation represented by the graph?

OpenStudy (acannell):

4:1? Is it like a ratio I need a clearer response.

OpenStudy (frostarmy):

no it is not a ratio.

OpenStudy (frostarmy):

I'm doing Sequences and Modeling with Functions

OpenStudy (acannell):

Oh...

OpenStudy (loser66):

Pick two special points on the graph (whose coordinates are integers), please

OpenStudy (frostarmy):

it doesn't give me any integers, just that right there

OpenStudy (loser66):

Why not? on the graph, you have (-2,2) and (2,0) Is it not that all of them integers?

OpenStudy (frostarmy):

well i just re-looked at it and i also see (6,-2 and -6,4)

OpenStudy (loser66):

Plenty of them there!! you can pick (-6,4) or (6,-2). They are special also

OpenStudy (loser66):

yup

OpenStudy (loser66):

Now, we need just 2 of them. I follow your points

OpenStudy (loser66):

What is slope-point formula?

OpenStudy (frostarmy):

y-y^1=m(x-x^1)

OpenStudy (frostarmy):

not sure if I'm right on that

OpenStudy (loser66):

not \(y^1\), it is \(y_1\)

OpenStudy (frostarmy):

yea that i don't know how to type it like that

OpenStudy (loser66):

So, we find m now. Ok?

OpenStudy (loser66):

\(m = ?\)

OpenStudy (frostarmy):

M would be the slope..right?

OpenStudy (loser66):

yup

OpenStudy (frostarmy):

ok let me see

OpenStudy (frostarmy):

so doing (-2,2) and (2,0) the slope would be -4/2 right?

OpenStudy (loser66):

\(m=\dfrac{\color{red}{y_2}-y_1}{\color{blue}{x_2}-x_1}\) Your points \((x_1,y_1)=(-2,2)\), and \((\color{blue}{x_2},\color{red}{y_2})=(\color{blue}{2},\color{red}{0})\) How can you have \(m =-\dfrac{4}{2}\) ?

OpenStudy (loser66):

redo, please

OpenStudy (frostarmy):

so -2/4 ?

OpenStudy (loser66):

yup, simplify to get -1/2, ok? Now, pick one point, take that m, plug all in formula of a line above \(y-y_1=m(x-x_1)\) and simplify

OpenStudy (frostarmy):

ok give me a min

OpenStudy (frostarmy):

so 0-2=-1/2 (2- (-2))